Kgabo completed 2/5 of a job. The next day he completed 3/8 of the remaining job. The part of the job that




is still left, given as a fraction, is


1
Expert's answer
2022-02-10T17:56:21-0500
"1-\dfrac{2}{5}-\dfrac{3}{8}(1-\dfrac{2}{5})=\dfrac{3}{8}"

The part of the job that is still left, given as a fraction, is "\dfrac{3}{8}."